OpenStudy (anonymous):

solve system by substution x=y x+2y=3

OpenStudy (anonymous):

If x=y, you can substitute x for y, or y for x to solve

OpenStudy (anonymous):

ok thanks alot

OpenStudy (anonymous):

If x=y, then x+2y=3 is the same as y+2y=3... therefore: 3y=3

OpenStudy (anonymous):

likewise: x+2y=3 is the same as x+2x=3, so: 3x=3